OpenStudy (anonymous):

What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths example plzzz last one yaaa

OpenStudy (ash2326):

http://en.wikipedia.org/wiki/Special_right_triangles Refer this if you get a doubt then reply here , I'll explain you:) @mak_12

OpenStudy (anonymous):

go ahead...

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

sorry wasnt by the computer but the question is up at the top @dpaInc

OpenStudy (anonymous):

@ash2326 so i looked at ur site n dont know what to put fo rhte answer im confuesed

OpenStudy (ash2326):

Hey mak_12 You know about a right triangle. It follows the Pythagoras theorem and one angle is right angle/

OpenStudy (ash2326):

@mak_12 ??

OpenStudy (anonymous):

right so whats specail ab it that it can be a right angle too?

OpenStudy (anonymous):

sorry computer terribly slow

OpenStudy (anonymous):

hey brb sorry step mom callin me 5 mins

OpenStudy (ash2326):

There can be infinite right triangles but there are a few right triangles where finding the sides is quite easy. Like a right triangle with the other two angles equal that's 45 each. So the other two sides will be equal and third side is hypotenuse.

OpenStudy (ash2326):

Or the right triangle with sides 3,4 5 where \(5^2=3^2+4^2\) or 8,15 and 17 where \({17}^2=8^2+{15}^2\) and so on. Here it's easy to find the sides. These are special right triangles

OpenStudy (anonymous):

thats a way batter to understand thak u soo much